A voltaic cell is constructed that uses the following reaction and operates at 298K:
Zn(s) + Ni
1
Expert's answer
2021-07-19T03:14:51-0400

"E_{cell}=E_{cell}^0=-\frac{RT}{nF}\ln\frac{P}{R}\\
E_{cell}=E_{cell}^0=-\frac{0.0591}{2}\ln\frac{0.137}{3.94}=0.55 V"